# Leveraging Large Language Models to Address Data Scarcity in Machine Learning: Applications in Graphene Synthesis

### v1

@@ 3 Results
We first evaluate the data imputation performance of all imputation methods by comparing the mean absolute error (MAE) of the imputed data against the true values, as detailed in section 2. The ground-truth values for each method are derived from the originally complete dataset before introducing artificial missing values. For each attribute, the values that were originally present in the mined dataset, are considered to be the known ground truth values for that attribute and used as references for both benchmarking and imputation steps. KNN imputation benchmark is performed by the statistical average of 50% known neighbors for the 50% artificial missing values. The imputed values for the artificial missing placeholders are compared to the ground-truth values which were removed in the previous step, to compute the deviation. The choice of 50% neighbors is done to ensure that KNN has a representative knowledge before imputation. For LLM-based methods, such as ChatGPT prompt imputation, there is no prior knowledge of neighbors involved. Instead, missing values are imputed based on the model’s pre-trained knowledge as well as the information provided through the prompt to guide the imputation.
@@
Pressure (mbar) exhibits a bimodal distribution in its raw form, and the imputed data from different methods generally follow this pattern. KNN imputation for pressure closely matches the raw data distribution (as shown in Figure 4). The imputation of pressure by GUIDE is slightly more diverse than KNN, with a greater population of imputed values near the first peak. However, MAPV is more inclined toward the second peak, where it concentrates most of the imputed values. This distributional behavior in MAPV is a result of the prompt instructions, which explicitly incorporate both peaks into the structure of the prompt and bias the imputation toward the second peak. On the other hand, CITEV exhibits a distribution similar to KNN, as it closely imitates the raw data distribution by populating most values around the first peak. The contextual information embedded in the prompt allows CITEV to replicate the underlying data structure more precisely, even placing fewer values between the peaks compared to KNN. A more diverse distribution that still aligns with the raw data is produced by the GIST method. GIST imputes pressure and other attributes using pre-defined language descriptors within GUIDE, making it more flexible in generating values in lower-population regions. As shown in Figure 3 GIST exhibits the lowest MAE for pressure imputation.
@@
To select the best classification model, we compare the performance of four different models – Decision Tree, Random Forest, XGBoost, and SVM – for binary classification on our scarce dataset using the GUIDE imputed dataset with a four dimensional substrate embedding featurization in the continuous variable space (GUIDE-SE(4)). As shown in Figure 6, tree-based models, including Random Forest, Decision Tree, and XGBoost, achieve perfect training scores but exhibit poor cross-validation performance. In contrast, the SVM classifier demonstrates convergence between training and cross-validation scores as the training size increases, highlighting its ability to better learn from scarce data and generalize to unseen instances. The tendency of tree-based models to memorize the training data stems from their higher complexity and expressivity. However, due to the limited data available, they are prone to overfitting, as indicated by the large gap between their training and cross-validation curves. Another key observation is that all models are constrained by the limitations of the scarce dataset. While SVM exhibits the least overfitting and achieves the highest validation accuracy of around 60%, the cross-validation scores of all models generally fall between 50% and 60%. This suggests that, regardless of the model architecture, the overall learning process is fundamentally restricted by data scarcity.
@@
To assess the impact of different substrate featurization methods, as detailed in Table 3, on learning performance, we compare the accuracy of an SVM binary classifier (monolayer vs. multilayer graphene) on GUIDE-imputed data across various featurization approaches. Figure 7 illustrates the classification accuracy for each method. The classifier performs similarly to a naive model (accuracy ¡ 0.5) when using label-encoded (LE) features. However, the LLM-standardized substrate features (SSF) without embeddings improve accuracy from 45% to 52% compared to LE (examples of SSF are shown in Supplementary Table 2). Additionally, applying the text-embedding-3-small model to the raw substrate attributes yields higher accuracy, particularly when using 4-dimensional (SE(4)) and 8-dimensional (SE(8)) embedding vectors. We further vectorize the key-value pairs of the standardized substrate features using the embedding model to evaluate accuracy, referred to as Standardized Features on Substrate Embedding (SFSE). The SFSE(4) method achieves the same accuracy as SE(8), but increasing the dimensions further to SFSE(8) results in lower accuracy. This decline is likely due to the trade-off between feature space dimensionality and data population. SE(8) effectively captures meaningful and contextual representations by leveraging the embedding model, extending into higher dimensions. However, for SFSE(8), the variability in input representation is lower (i.e., due to standardization), leading to sparse contextual embeddings that primarily reflect the predefined contexts in the key-value pairs. Classification performance metrics for different featurization methods are detailed in Supplementary Table 3.
@@
We also investigated the impact of discretizing the feature space on classification performance, as discretization can be beneficial in data-scarce scenarios. Figure 7 compares the accuracy of the SVM binary classifier for GUIDE-imputed data with the SE(4) substrate featurization across different discretization approaches, as detailed in Table 4. K-means binning with six clusters (K-Means(6)) slightly reduces accuracy compared to SE(4) in the continuous space. This decline is likely due to suboptimal bin splits, as K-means minimizes intracluster variance, leading to non-uniform bin divisions. In contrast, equal-width binning with ten bins (Eq-Width(10)) yields higher accuracy than both K-Means(6) and continuous SE(4). Eq-Width(10) divides attributes into ten equally spaced bins, ensuring uniform discretization by definition, which results in improved accuracy over K-means-based binning. Conversely, equal-frequency binning with twelve bins (Eq-Freq(12)) leads to lower accuracy, likely due to the same issue of suboptimal or nonuniform bin splits. The effect of different bin sizes is discussed in Supplementary Figure 9. . Additionally, we examine a supervised discretization approach using decision trees (DT), which optimizes bin splits by minimizing the Gini impurity, ensuring bins are formed based on attribute-target correlations. This method is particularly effective, as it clusters bins in a more informed manner than unsupervised binning methods. The accuracy of an 80-20 train-test split with DT (DT(80-20)) is significantly higher than that of unsupervised binning methods and is further improved with a 90-10 split (DT(90-10)). The classification performance metrics for different discretization methods are reported in Supplementary Table 4.
@@
Finally, we evaluate the cumulative impact of all the strategies presented in this study for improving classification performance on scarce, inhomogeneous data. To do so, we compare the most basic data processing and feature engineering approach against the most effective strategy identified in this study for the SVM classifier. The least processed approach consists of KNN-imputed data with label-encoded substrate featurization in the continuous feature space (KNN-LabelEncoded-Continuous). In contrast, the most advanced method utilizes GIST-imputed data with text-embedding-based substrate featurization (8-dimensional) and decision tree discretization (GIST-SE(8)-DT). The exception to this is for the ternary case where the GUIDE-imputed data with 4 dimensional substrate embedding and decision tree discretization, shows the highest accuracy (GUIDE-SE(4)-DT). Figure 9 illustrates t-distributed stochastic neighbor embedding (t-SNE) maps for the binary SVM classifier, comparing the worst-performing (KNN-LabelEncoded-Continuous) and best-performing (GIST-SE(8)-DT) models. For the ternary case it shows the comparison between the worst-performing (KNN-LabelEncoded-Continuous) and best performing (GUIDE-SE(4)-DT). These results show a significant accuracy improvement from 39% to 65% for binary classification (Figure 9(a)-(b)) and from 52% to 72% for ternary classification (Figure 9(c)-(d)).
@@
Additionally, we fine-tune two ChatGPT-4o models to directly classify the binary and ternary classes and compare their performance with the SVM classifier. The same training datasets were used to fine-tune the models. The technical details are presented in Supplementary Note 5. The predictions of both fine-tuned ChatGPT-4o and ChatGPT-4o-mini are shown in Figure 10 along with the SVM classifier. The LLM-based classification is significantly outperformed by the more standard SVM classifier for both binary and ternary cases. The scores reflect the performance of both approaches under scarce data conditions, showing that traditional ML models combined with LLM-based data engineering techniques outperform fine-tuned LLM models alone. Detailed accuracy reports are provided in Supplementary Table 5 and the cross-validation curve for fine-tuning process is shown in Supplementary Figure 13. The potential of an LLM fine-tuned on larger datasets remains speculative and is beyond the scope of this study. Our general guidelines can leverage LLM to improve conventional data engineering strategies to optimize learning for numerical models.
@@ Acknowledgment
The authors declare that they have no financial or non-financial competing interests.
@@
The raw dataset contains data entries for attributes corresponding to the graphene growth conditions and graphene number of layers, i.e., the target variable. The raw attributes are Pressure (mbar), Temperature (°C), Growth Time (min), Substrate, Number of Graphene Layers, and the gas Flow Rates (sccm). These attributes and their corresponding entries are mined from the experimental chemical vapor deposition (CVD) graphene growth literature 57, 58, 59, 60, 50, 61, 62, 63, 64. The gas flow rates are initially extracted as a dictionary- {gas: flow rate}, later separated into different columns. Out of many gases reported in the literature, we select the top 5 most common gases – H2, CH4, C2H4, Ar, and C2H2. This is to keep the number of attributes to the minimum, which is especially important to address the curse of dimensionality for scarce datasets. The Suppelementary File “DataProcessing.ipynb” shows the count of each attribute before the imputation process and can be used to reproduce the results for the imputation step. The collected data are stored in the “graphene_growth_conditions_layers.csv” file.
@@ 1. GUIDE:
System: You are a CVD expert with access to experimental and/or simulated data from literature, which you can understand to derive correlations between known and unknown quantities for given datasets.
@@
Pressure: {context_vars[’pressure’]} mbar, Temperature: {context_vars[’temperature’]} °C, Growth Time: {context_vars[’growth_time’]} min, Substrate: {context_vars[’substrate’]}, Gas Flow Rates: H2: {context_vars[’h2’]}, C2H4: {context_vars[’c2h4’]}, Ar: {context_vars[’ar’]}, C2H2: {context_vars[’c2h2’]}. Can you determine the missing value of {target_variable}? Please provide only the missing value as a single number, without any additional text.
@@ 2. MAP:
System: You are a CVD expert with access to experimental and/or simulated data from literature, which you can understand to derive correlations between known and unknown quantities for given datasets. Now, I would like you to help me impute the missing value for {target_variable}. Please use your scientific knowledge to impute this value, based only on the current row’s context.
@@
  • H2: Highly right-skewed, with most values near 0 and a few extending to higher values.
  • CH4: Right-skewed, with the majority of the values concentrated near 0.
  • Ar: Very right-skewed, with most values concentrated near 0 and very few extending up to 10,000.
@@ 6. GIST:
The analysis is carried out with normalized attributes to compensate for the scaling factor and different orders of magnitude of the attributes, as shown in Supplementary Figure 4. We use a standard scaler to normalize the ground truth values between 0 and 1. Some imputation values fall above the normalized scale because they are plotted with respect to the ground truth normalized values. When scaling with the min-max scaler, we need to select the minimum and maximum values of either the ground-truth or the imputed value range. This can lead to some of the imputed values to fall out of the scale which simply means that the imputed values are greater than the maximum ground-truth value for the corresponding attribute. The red, blue, and yellow points represent the ground truth, averaged imputes, and minimum imputed values, respectively. The averaged imputed values are computed over six iterations, providing an overall est. In contrast, the minimum imputed values correspond to the instance with the lowest deviation from the ground truth, highlighting the most accurate predictions. The imputed values for 6 attributes with the GIST prompt are shown in Supplementary Figure 4.
@@
We choose K=5 for the KNN imputation process which accounts for a reasonable number of neighbors for the scarce dataset. Higher K values can bias the predictions while lower K values might not capture the statistical variance. Moreover, $K=4$ and $K=6$ follows the same distribution as K=5 as shown in Supplementary Table 1 with the average MAE and MSE values (deviations from ground-truth) for the imputation of all attributes, allowing us to limit the value to $K=5$.
@@
Supplementary Figure 5 shows the imputation analysis for KNN. We use 50 % of the ground truth data for KNN benchmarking while keeping the remaining 50 % as validation set. The ChatGPT imputation is carried on without any prior knowledge of the ground truth data while the KNN requires some statistical distribution to work on. A higher split of training data would leave just a few points for validation, so we stick to 50 % as a reasonable split. For the imputation step post benchamrking, we use all the known ground truth values. Supplementary Figure 5 shows the KNN imputation benchmarking.
@@
The featurization of the substrate attribute involves addressing diverse categories and absence of a standard format. Key techniques used include:
@@
Note: We use the continuous-valued attributes to test the featurization techniques and pick the best choice. This would be important to keep the featurization step different than the discretizing step which would require its own testing before combining with the featurization techniques.
@@
An important step before training with the features is to reduce the dimensions of the embedding vectors to avoid the dimension problem. This can be manually done by principal component analysis but we leverage OpenAI’s inbuilt method to reduce the dimensions to 4, 8 or 16 as these are within the dimension limit for the small dataset. Supplementary Figure 7 shows the code snippet for the embedding method incorporated for this case.
@@
The featurization techniques are assessed by an SVM model, the details of which is are given in Supplementary Note 4.2. As shown in Supplementary Table 3, the classification performance is highest for the substrate embedding method with 8 reduced dimensions (SE-8), showing a substantial increase over the label encoded method.
@@
We discretize the variable space to address attributes with heavily skewed distributions, such as Pressure, Growth Time, and Gas Flow Rates. Training models over continuous values for these attributes often undermines both accuracy and usability, as the model may fail to capture meaningful patterns in the data. By converting continuous values into categorical labels, discretization simplifies the data and enhances model performance and interpretability. In this study, we examines the following types of Discretization Techniques:
@@
To discretize without increasing the dimensions of the dataset, we use label encoding. By assigning integer labels to the bins created through the selected discretization technique, we keep the transformed data compact while maintaining its interpretability.
1. Equal-width Binning: Divides the range into bins of equal width.

### v2

@@ 3 Results
We first evaluate the data imputation performance of all imputation methods by comparing the mean absolute error (MAE) of the imputed data against the true values, as detailed in section 2. The ground-truth values for each method are derived from the originally complete dataset before introducing artificial missing values. For each attribute, the values that were originally present in the mined dataset, are considered to be the known ground truth values for that attribute and used as references for both benchmarking and imputation steps. KNN imputation benchmark is performed by the statistical average of 50% known neighbors for the 50% artificial missing values. The imputed values for the artificial missing placeholders are compared to the ground-truth values, which were removed in the previous step, to compute the deviation. The choice of 50% neighbors is done to ensure that KNN has representative knowledge before imputation. For LLM-based methods, such as ChatGPT prompt imputation, there is no prior knowledge of neighbors involved. Instead, missing values are imputed based on the model’s pre-trained knowledge as well as the information provided through the prompt to guide the imputation.
@@
Pressure (mbar) exhibits a bimodal distribution in its raw form, and the imputed data from different methods generally follow this pattern. KNN imputation for pressure closely matches the raw data distribution (as shown in Figure 4). The imputation of pressure by GUIDE is slightly more diverse than KNN, with a greater population of imputed values near the first peak. However, MAPV is more inclined toward the second peak, where it concentrates most of the imputed values. This distributional behavior in MAPV is a result of the prompt instructions, which explicitly incorporate both peaks into the structure of the prompt and bias the imputation toward the second peak. On the other hand, CITEV exhibits a distribution similar to KNN, as it closely imitates the raw data distribution by populating most values around the first peak. The contextual information embedded in the prompt allows CITEV to replicate the underlying data structure more precisely, even placing fewer values between the peaks compared to KNN. A more diverse distribution that still aligns with the raw data is produced by the GIST method. GIST imputes pressure and other attributes using pre-defined language descriptors within GUIDE, making it more flexible in generating values in lower-population regions. As shown in Figure 3, GIST exhibits the lowest MAE for pressure imputation.
@@
To select the best classification model, we compare the performance of four different models – Decision Tree, Random Forest, XGBoost, and SVM – for binary classification on our scarce dataset using the GUIDE imputed dataset with a four-dimensional substrate embedding featurization in the continuous variable space (GUIDE-SE(4)). As shown in Figure 6, tree-based models, including Random Forest, Decision Tree, and XGBoost, achieve perfect training scores but exhibit poor cross-validation performance. In contrast, the SVM classifier demonstrates convergence between training and cross-validation scores as the training size increases, highlighting its ability to better learn from scarce data and generalize to unseen instances. The tendency of tree-based models to memorize the training data stems from their higher complexity and expressivity. However, due to the limited data available, they are prone to overfitting, as indicated by the large gap between their training and cross-validation curves. Another key observation is that all models are constrained by the limitations of the scarce dataset. While SVM exhibits the least overfitting and achieves the highest validation accuracy of around 60%, the cross-validation scores of all models generally fall between 50% and 60%. This suggests that, regardless of the model architecture, the overall learning process is fundamentally restricted by data scarcity.
@@
To assess the impact of different substrate featurization methods, as detailed in Table 3, on learning performance, we compare the accuracy of an SVM binary classifier (monolayer vs. multilayer graphene) on GUIDE-imputed data across various featurization approaches. Figure 7 illustrates the classification accuracy for each method. The classifier performs similarly to a naive model (accuracy ¡ 0.5) when using label-encoded (LE) features. However, the LLM-standardized substrate features (SSF) without embeddings improve accuracy from 45% to 52% compared to LE (examples of SSF are shown in Supplementary Table 2). Additionally, applying the text-embedding-3-small model to the raw substrate attributes yields higher accuracy, particularly when using 4-dimensional (SE(4)) and 8-dimensional (SE(8)) embedding vectors. We further vectorize the key-value pairs of the standardized substrate features using the embedding model to evaluate the accuracy, referred to as Standardized Features on Substrate Embedding (SFSE). The SFSE(4) method achieves the same accuracy as SE(8), but increasing the dimensions further to SFSE(8) results in lower accuracy. This decline is likely due to the trade-off between feature space dimensionality and data population. SE(8) effectively captures meaningful and contextual representations by leveraging the embedding model, extending into higher dimensions. However, for SFSE(8), the variability in input representation is lower (i.e., due to standardization), leading to sparse contextual embeddings that primarily reflect the predefined contexts in the key-value pairs. Classification performance metrics for different featurization methods are detailed in Supplementary Table 3.
@@
We also investigated the impact of discretizing the feature space on classification performance, as discretization can be beneficial in data-scarce scenarios. Figure 7 compares the accuracy of the SVM binary classifier for GUIDE-imputed data with the SE(4) substrate featurization across different discretization approaches, as detailed in Table 4. K-means binning with six clusters (K-Means(6)) slightly reduces accuracy compared to SE(4) in the continuous space. This decline is likely due to suboptimal bin splits, as K-means minimizes intracluster variance, leading to non-uniform bin divisions. In contrast, equal-width binning with ten bins (Eq-Width(10)) yields higher accuracy than both K-Means(6) and continuous SE(4). Eq-Width(10) divides attributes into ten equally spaced bins, ensuring uniform discretization by definition, which results in improved accuracy over K-means-based binning. Conversely, equal-frequency binning with twelve bins (Eq-Freq(12)) leads to lower accuracy, likely due to the same issue of suboptimal or nonuniform bin splits. The effect of different bin sizes is discussed in Supplementary Figure 9. Additionally, we examine a supervised discretization approach using decision trees (DT), which optimizes bin splits by minimizing the Gini impurity, ensuring bins are formed based on attribute-target correlations. This method is particularly effective, as it clusters bins in a more informed manner than unsupervised binning methods. The accuracy of an 80-20 train-test split with DT (DT(80-20)) is significantly higher than that of unsupervised binning methods and is further improved with a 90-10 split (DT(90-10)). The classification performance metrics for different discretization methods are reported in Supplementary Table 4.
@@
Finally, we evaluate the cumulative impact of all the strategies presented in this study for improving classification performance on scarce, inhomogeneous data. To do so, we compare the most basic data processing and feature engineering approach against the most effective strategy identified in this study for the SVM classifier. The least processed approach consists of KNN-imputed data with label-encoded substrate featurization in the continuous feature space (KNN-LabelEncoded-Continuous). In contrast, the most advanced method utilizes GIST-imputed data with text-embedding-based substrate featurization (8-dimensional) and decision tree discretization (GIST-SE(8)-DT). The exception to this is for the ternary case where the GUIDE-imputed data with four-dimensional substrate embedding and decision tree discretization shows the highest accuracy (GUIDE-SE(4)-DT). Figure 9 illustrates t-distributed stochastic neighbor embedding (t-SNE) maps for the binary SVM classifier, comparing the worst-performing (KNN-LabelEncoded-Continuous) and best-performing (GIST-SE(8)-DT) models. For the ternary case, it shows the comparison between the worst-performing (KNN-LabelEncoded-Continuous) and the best-performing (GUIDE-SE(4)-DT). These results show a significant accuracy improvement from 39% to 65% for binary classification (Figure 9(a)-(b)) and from 52% to 72% for ternary classification (Figure 9(c)-(d)).
@@
Additionally, we fine-tune two ChatGPT-4o models to directly classify the binary and ternary classes and compare their performance with the SVM classifier. The same training datasets were used to fine-tune the models. The technical details are presented in Supplementary Note 5. The predictions of both fine-tuned ChatGPT-4o and ChatGPT-4o-mini are shown in Figure 10 along with the SVM classifier. The LLM-based classification is significantly outperformed by the more standard SVM classifier for both binary and ternary cases. The scores reflect the performance of both approaches under scarce data conditions, showing that traditional ML models combined with LLM-based data engineering techniques outperform fine-tuned LLM models alone. Detailed accuracy reports are provided in Supplementary Table 5, and the cross-validation curve for fine-tuning process is shown in Supplementary Figure 13. The potential of an LLM fine-tuned on larger datasets remains speculative and is beyond the scope of this study. Our general guidelines can leverage LLM to improve conventional data engineering strategies to optimize learning for numerical models.
@@ Acknowledgment
The authors declare that they have no competing financial or non-financial interests.
@@
The raw dataset contains data entries for attributes corresponding to the graphene growth conditions and graphene number of layers, i.e., the target variable. The raw attributes are Pressure (mbar), Temperature (°C), Growth Time (min), Substrate, Number of Graphene Layers, and gas Flow Rates (sccm). These attributes and their corresponding entries are mined from the experimental chemical vapor deposition (CVD) graphene growth literature 57, 58, 59, 60, 50, 61, 62, 63, 64. The gas flow rates are initially extracted as a dictionary- {gas: flow rate}, later separated into different columns. Out of many gases reported in the literature, we select the top 5 most common gases – H2, CH4, C2H4, Ar, and C2H2. This is to keep the number of attributes to a minimum, which is especially important to address the curse of dimensionality for scarce datasets. The Supplementary File “DataProcessing.ipynb” shows the count of each attribute before the imputation process and can be used to reproduce the results for the imputation step. The collected data are stored in the “graphene_growth_conditions_layers.csv” file.
@@ 1. GUIDE:
System: You are a CVD expert with access to experimental and/or simulated data from the literature, which you can understand to derive correlations between known and unknown quantities for given datasets.
@@
Pressure: {context_vars[’pressure’]} mbar, Temperature: {context_vars[’temperature’]} °C, Growth Time: {context_vars[’growth_time’]} min, Substrate: {context_vars[’substrate’]}, Gas Flow Rates: H2: {context_vars[’h2’]}, C2H4: {context_vars[’c2h4’]}, Ar: {context_vars[’ar’]}, C2H2: {context_vars[’c2h2’]}. Can you determine the missing value of {target_variable}? Please provide only the missing value as a single number without any additional text.
@@ 2. MAP:
System: You are a CVD expert with access to experimental and/or simulated data from literature, which you can understand to derive correlations between known and unknown quantities for given datasets. Now, I would like you to help me impute the missing value for {target_variable}. Please use your scientific knowledge to impute this value based only on the current row’s context.
@@
  • H2: Highly right-skewed, with most values near zero and a few extending to higher values.
  • CH4: Right-skewed, with most values concentrated near 0.
  • Ar: Very right-skewed, with most values concentrated near zero and very few extending up to 10,000.
@@ 5. CITEV:
Here’s the current attribute data, you need to impute based on this knowledge and your knowledge of the literature: Also make sure to impute values around the modes and also in regions of scarely populated data, to have a diverse data distribution {data_preprocessed[target_variable].dropna().iloc[:len(data_preprocessed[target_variable].dropna()) // 2].to_string()
@@ 6. GIST:
The analysis is carried out with normalized attributes to compensate for the scaling factor and different orders of magnitude of the attributes, as shown in Supplementary Figure 4. We use a standard scaler to normalize the ground truth values between 0 and 1. Some imputation values fall above the normalized scale because they are plotted with respect to the ground truth normalized values. When scaling with the min-max scaler, we need to select the minimum and maximum values of either the ground truth or the imputed value range. This can lead to some of the imputed values falling out of the scale, which simply means that the imputed values are greater than the maximum ground-truth value for the corresponding attribute. The red, blue, and yellow points represent the ground truth, averaged imputes, and minimum imputed values, respectively. The averaged imputed values are computed over six iterations, providing an overall est. In contrast, the minimum imputed values correspond to the instance with the lowest deviation from the ground truth, highlighting the most accurate predictions. The imputed values for six attributes with the GIST prompt are shown in Supplementary Figure 4.
@@
We chose K=5 for the KNN imputation process, which accounts for a reasonable number of neighbors for the scarce dataset. Higher K values can bias the predictions, while lower K values might not capture the statistical variance. Moreover, $K=4$ and $K=6$ follow the same distribution as K=5 as shown in Supplementary Table 1 with the average MAE and MSE values (deviations from ground-truth) for the imputation of all attributes, allowing us to limit the value to $K=5$.
@@
Supplementary Figure 5 shows the imputation analysis for KNN. We use 50 % of the ground truth data for KNN benchmarking while keeping the remaining 50 % as a validation set. The ChatGPT imputation is carried on without any prior knowledge of the ground truth data, while the KNN requires some statistical distribution to work on. A higher split of training data would leave just a few points for validation, so we stick to 50 % as a reasonable split. For the imputation step post benchmarking, we use all the known ground truth values. Supplementary Figure 5 shows the KNN imputation benchmarking.
@@
The featurization of the substrate attribute involves addressing diverse categories and the absence of a standard format. Key techniques used include:
@@
Note: We use the continuous-valued attributes to test the featurization techniques and pick the best choice. It would be important to keep the featurization step different from the discretizing step, which would require its own testing before being combined with the featurization techniques.
@@
An important step before training with the features is to reduce the dimensions of the embedding vectors to avoid the dimension problem. This can be manually done by principal component analysis, but we leverage OpenAI’s inbuilt method to reduce the dimensions to 4, 8, or 16 as these are within the dimension limit for the small dataset. Supplementary Figure 7 shows the code snippet for the embedding method incorporated for this case.
@@
The featurization techniques are assessed by an SVM model, the details of which are given in Supplementary Note 4.2. As shown in Supplementary Table 3, the classification performance is highest for the substrate embedding method with eight reduced dimensions (SE-8), showing a substantial increase over the label encoded method.
@@
We discretize the variable space to address attributes with heavily skewed distributions, such as Pressure, Growth Time, and Gas Flow Rates. Training models over continuous values for these attributes often undermines accuracy and usability, as the model may fail to capture meaningful patterns in the data. Discretization simplifies the data and enhances model performance and interpretability by converting continuous values into categorical labels. In this study, we examine the following types of Discretization Techniques:
@@
We use label encoding to discretize without increasing the dataset’s dimensions. By assigning integer labels to the bins created through the selected discretization technique, we keep the transformed data compact while maintaining its interpretability.
1. Equal-width Binning: Divide the range into bins of equal width.